OpenStudy (anonymous):

If 3/2 divided by 1/4 = n, then n is between?

OpenStudy (anonymous):

Forget between, you can calculate n exactly. (Pro-tip: dividing by a number is the same as multiplying by the reciprocal).

OpenStudy (anonymous):

Well I actually need to know between specifically. This is part of an accuplacer I'm studying for

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{2} \div \cfrac{1}{4}\implies \cfrac{\frac{3}{2}}{\frac{1}{4}}\quad recall\implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\quad thus \\ \quad \\ \cfrac{\frac{3}{2}}{\frac{1}{4}}\implies \cfrac{3}{2}\times \cfrac{4}{1}\)

OpenStudy (anonymous):

Thanks to both of you, you're both really helpful

OpenStudy (jdoe0001):

as walking_stick said, find "n" once you find "n"... then you should be able to tell who's LOWER or HIGHER than "n" and thus in between who it's in

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

:)